Presented below is information related to Splish Company.

Date

Ending Inventory
(End-of-Year Prices)

Price
Index

December 31, 2017

$ 83,600 100

December 31, 2018

261,723 231

December 31, 2019

258,456 264

December 31, 2020

292,006 286

December 31, 2021

349,580 308

December 31, 2022

421,399 319


Compute the ending inventory for Splish Company for 2017 through 2022 using the dollar-value LIFO method.

Ending Inventory

2017

$enter a dollar amount

2018

$enter a dollar amount

2019

$enter a dollar amount

2020

$enter a dollar amount

2021

$enter a dollar amount

2022

$enter a dollar amount
Answer #1


Compute the ending inventory for Splish Company for 2017 through 2022 using the dollar-value LIFO method.

Ending Inventory

2017

83600

2018

261723/2.31 = 113300 (83600+29700*2.31) = 152207

2019

258456/2.64 = 97900 (83600+14300*2.31) = 116633

2020

292006/2.86 = 102100 (116633+4200*2.86) = 128645

2021

349580/3.08 = 113500 (128645+11400*3.08) = 163757

2022

421399/3.19 = 132100 (163757+18600*3.19) = 223091